# MS-0001-1922.65

### v2

@@
As generative artificial intelligence (AI) tools become widely adopted, large language models (LLMs) are increasingly involved on both sides of decision-making processes, ranging from hiring to content moderation. This dual adoption raises a critical question: do LLMs systematically favor content that resembles their own outputs? Prior research in computer science has identified self-preference bias—the tendency of LLMs to favor their own generated content—but its real-world implications have not been empirically evaluated. We focus on the hiring context, where job applicants often rely on LLMs to refine resumes, while employers deploy them to screen those same resumes. Using a large-scale controlled resume correspondence experiment, we find that LLMs consistently prefer resumes generated by themselves over those written by humans or produced by alternative models, even when content quality is controlled. The bias against human-written resumes is particularly substantial, with self-preference bias ranging from $68\%$ to $88\%$ across major commercial and open-source models. To assess labor market impact, we simulate realistic hiring pipelines across 24 occupations. These simulations show that candidates using the same LLM as the evaluator are $23\%$ to $60\%$ more likely to be shortlisted than equally qualified applicants submitting human-written resumes, with the largest disadvantages observed in business-related fields such as sales and accounting. We further demonstrate that this bias can be reduced by more than $50\%$ through simple interventions targeting LLMs’ self-recognition capabilities. These findings highlight an emerging but previously overlooked risk in AI-assisted decision making and call for expanded frameworks of AI fairness that address not only demographic-based disparities, but also biases in AI-AI interactions.
@@ 1 Introduction
The rapid development and commercialization of generative artificial intelligence (AI) have made large language models (LLMs) widely accessible across both professional and everyday contexts. As these tools become embedded in diverse workflows, they are increasingly involved on both sides of content generation and evaluation. In hiring, for instance, applicants often use LLMs to draft or refine their resumes, while employers leverage similar tools to screen or rank candidates (New York State Society of CPAs 2024, ResumeBuilder.com 2023, Wiles et al. 2025). On social media, users may rely on LLMs to help compose or polish posts, while platforms employ LLMs to moderate content, e.g., flagging, categorizing, or filtering user submissions (Kumar et al. 2024). In academia, researchers use LLMs to improve their manuscripts, while conferences and journals are beginning to experiment with LLM-assisted peer review (Thakkar et al. 2025, Naddaf 2025). Similar patterns are emerging in education and customer service, where LLMs support both communication and assessment tasks. In these domains, LLMs now function as both content generators and evaluators, giving rise to a new class of AI–AI interactions with significant implications for human decision-making and outcomes (Laurito et al. 2025).
@@
Recent research in computer science has identified a behavioral tendency in LLMs known as self-preference, i.e., the inclination of a model to favor content it generated itself over that written by humans or produced by alternative models (Panickssery et al. 2024). While this phenomenon has been documented in benchmark evaluations (Zheng et al. 2023, Bai et al. 2023), its implications for real-world decision-making remain understudied. As LLMs are deployed in high-stakes settings where they may evaluate content also generated by LLMs, self-preference introduces a novel form of bias. Unlike traditional biases rooted in demographic disparities (Sheng et al. 2019), this bias emerges from AI-AI interactions, in which the model’s own evaluative behavior favors outputs aligned with its generative patterns.
To date, research on AI fairness has primarily focused on protected attributes such as race, gender, and age (Abid et al. 2021, Zhao et al. 2024, Liu et al. 2024). A growing body of evidence shows that LLMs can reproduce and even amplify social stereotypes along these lines, often due to biased training data or inadequate debiasing interventions (Nangia et al. 2020, Nadeem et al. 2021, Kotek et al. 2023, Cheng et al. 2023). As a result, regulatory frameworks and algorithmic audits have largely concentrated on mitigating harms associated with these well-defined dimensions of demographic attributes (Mickel 2024, Wright et al. 2024).
@@
In contrast, self-preference bias reflects a different and underexplored phenomenon that emerges from AI–AI interactions. Rather than being tied to demographic attributes, it depends on the source of the content under evaluation—specifically, whether it was produced by the evaluating LLM model itself. This bias is inherently interactional: it arises when LLMs are asked to judge content that may share stylistic or linguistic patterns with their own generative outputs. As such, it poses a new challenge for AI fairness, one that is not addressed by existing safeguards focused on demographic disparities. If left unchecked, self-preference could subtly distort evaluative processes across hiring, education, publishing, and more—privileging those who employ the same AI system used for evaluation (i.e., the “right” tool from the model’s perspective) while disadvantaging those who use different tools or none at all. Addressing this issue will require expanding current fairness frameworks to account for interactions between LLM models and their potential to shape outcomes in an increasingly AI-mediated world.
@@
Among the many domains where self-preference bias may arise, algorithmic hiring is particularly consequential. Employers are increasingly adopting LLMs to streamline resume screening and candidate ranking, often as part of automated workflows that support human decision-making (Gan et al. 2024, Kim et al. 2024, ResumeBuilder 2024, Sarumathi et al. 2025, Wiles and Horton 2025). Unlike traditional keyword-matching systems, LLMs can evaluate resumes in a more holistic manner—synthesizing content, inferring intent, and making contextual judgments beyond simple heuristics (Pritchett 2025). While this shift promises greater efficiency and scalability, it also magnifies the potential for bias. If an LLM systematically favors resumes that reflect its own generative style, it may confer unwarranted advantages to applicants who happen to use the same model to compose their materials. In such cases, evaluations may be influenced less by the substantive quality of a candidate’s credentials and more by superficial stylistic alignment with the evaluator LLM. In effect, this bias rewards access to specific generative technologies and penalizes those without it, even when applicants are otherwise equally qualified.
@@
In this paper, we provide the first empirical evidence that self-preference bias can distort candidate evaluations in algorithmic hiring. Specifically, we examine whether LLMs, when deployed as evaluators, systematically favor resumes they generated themselves over otherwise equivalent resumes written by humans or produced by alternative models. To test this, we construct a large-scale resume correspondence experiment using a real-world dataset of 2,245 human-written resumes, sourced from a professional resume-building platform prior to the widespread adoption of generative AI. For each resume, we generate multiple counterfactual versions using a range of state-of-the-art LLMs, including GPT-4o, GPT-4o-mini, GPT-4-turbo, LLaMA 3.3-70B, Mistral-7B, Qwen 2.5-72B, and Deepseek-V3. Having content quality controlled, we assess whether these LLMs exhibit systematic bias in favor of their own outputs when acting as evaluators.
@@
We distinguish between two forms of self-preference bias: LLM-vs-Human, where a model prefers its own generated content over a human-written equivalent; and LLM-vs-LLM, where a model favors its own output over content produced by a different LLM. We find strong and consistent evidence of LLM-vs-Human self-preferencing across most models. Larger systems—such as GPT-4o, GPT-4-turbo, DeepSeek-V3, Qwen-2.5-72B, and LLaMA 3.3-70B—exhibit particularly strong bias, exceeding $68\%$ even after controlling for content quality and reaching over $80\%$ for GPT-4o, Qwen-2.5-72B, and LLaMA 3.3-70B. By contrast, LLM-vs-LLM self-preferencing is more heterogeneous. DeepSeek-V3 shows the strongest bias in this setting, preferring its own outputs by $69\%$ against LLaMA 3.3-70B and $28\%$ against GPT-4o. GPT-4o and LLaMA 3.3-70B, in comparison, do not display consistent preferences when evaluating content generated by other models.
@@
To mitigate AI self-preference, we propose two simple yet effective strategies that directly target the underlying mechanism of self-recognition—a model’s ability to implicitly identify content it generated. The first strategy uses system prompting to explicitly instruct models to ignore the origin of resumes and focus only on substantive content. The second strategy employs a majority voting ensemble, combining the evaluator model with smaller models that exhibit weaker self-recognition, thereby diluting the bias of any single LLM. Across all tested LLMs, these interventions reduce LLM-vs-Human self-preference by more than $60\%$, with GPT-4o’s bias falling from over $90\%$ to below $50\%$. These results demonstrate that while self-preference bias is widespread and consequential, it is not immutable: straightforward design interventions can substantially improve fairness in LLM-based hiring evaluations.
Together, these findings contribute to a deeper understanding of how generative AI tools can unintentionally reinforce inequities in algorithmic evaluation. We document and quantify self-preference bias in the context of resume screening by developing a measurement framework grounded in established fairness metrics. In addition, we show that self-preference bias can be substantially reduced through targeted interventions informed by LLM’s self-recognition behavior. In doing so, the study introduces a novel and practical perspective on AI fairness—one that moves beyond concerns about demographic disparities to address interactional biases that arise when AI systems evaluate content they themselves could have produced.
@@ 2 Literature Review
Our research contributes to three streams of literature.
2.1 Fairness and Bias in Algorithmic Hiring
@@
Discrimination in labor markets has long been a central concern for both policymakers and economists, with extensive research documenting that certain race, gender, and age subpopulations face unfair treatment from recruiters in the hiring process (Bertrand and Mullainathan 2004, Kline et al. 2022, Neumark 2018). Concerningly, studies such as Datta et al. (2015) and Lambrecht and Tucker (2019) have further shown that, beyond human bias, algorithms deployed in hiring systems can also encode discriminatory behavior that disadvantages particular demographic groups. In recent years, the rise of LLMs with their increasing adoption in hiring systems has prompted researchers to empirically evaluate whether emerging LLM-powered hiring systems perpetuate similar forms of algorithmic discrimination. For instance, Veldanda et al. (2023) replicated Bertrand and Mullainathan’s correspondence experiment to investigate LLM-driven hiring bias across gender, race, maternity status, pregnancy status, and political affiliation. Additionally, An et al. (2024) explored whether LLMs exhibit race- and gender-based discrimination through an experiment, in which LLMs write an email to a named job candidate about a hiring decision. Likewise, Nghiem et al. (2024) analyzed the extent to which LLMs exhibit bias toward applicants based on their first names when making employment recommendations.
Our paper contributes to this rising literature by examining an overlooked bias in algorithmic hiring—the self-preference bias where LLMs disproportionately favor their own model-generated outputs over human or alternative LLM-generated responses. By uncovering this novel bias, our work highlights an emergent risk in algorithmic hiring systems that has yet to be addressed in current regulatory frameworks.
2.2 Self-Preference Bias in LLM-as-a-Judge

### v3

# AI Self-preferencing in Algorithmic Hiring: Empirical Evidence and Insights

@@
Problem definition: As artificial intelligence (AI) tools become widely adopted, large language models (LLMs) are increasingly involved on both sides of decision-making processes, ranging from hiring to content moderation. This dual adoption raises a critical question: do LLMs systematically favor content that resembles their own outputs? Prior research in computer science has identified self-preference bias, i.e., the tendency of LLMs to favor their own generated content, but its real-world implications have not been empirically evaluated. Methodology/results: We examine this question in the context of algorithmic hiring, a high-volume screening process central to workplace operations. In this setting, job applicants increasingly rely on LLMs to refine resumes, while employers deploy similar tools to screen those same materials. Using a large-scale controlled resume correspondence experiment, we find that LLMs consistently prefer resumes generated by themselves over those written by humans or produced by alternative models, even when content quality is controlled. The bias against human-written resumes is particularly substantial, with self-preference bias ranging from $67\%$ to $82\%$ across major commercial and open-source models. Managerial implications: To quantify the operational impact of self-preference bias, we simulate realistic hiring pipelines across $24$ occupations. These simulations show that candidates using the same LLM as the evaluator are $23\%$ to $60\%$ more likely to be shortlisted than equally qualified applicants submitting human-written resumes, with the largest disadvantages observed in business-related fields such as sales and accounting. We further demonstrate that, in many cases, this bias can be reduced by more than $50\%$ through simple interventions that target LLMs’ self-recognition capabilities. These findings highlight an emerging but previously overlooked risk in AI-assisted decision making and call for expanded frameworks of AI fairness in business operations that address biases from AI-AI interactions.
@@ 1 Introduction
The rapid development and commercialization of artificial intelligence (AI) have made large language models (LLMs) widely accessible across both professional and everyday contexts. As these tools become embedded in diverse workflows, they are increasingly involved on both sides of content generation and evaluation. In hiring, for instance, applicants often use LLMs to draft or refine their resumes, while employers leverage similar tools to screen or rank candidates (New York State Society of CPAs 2024, ResumeBuilder.com 2023, Wiles et al. 2025). On social media, users frequently use LLMs to help compose or polish posts, while platforms may employ LLMs to moderate content, e.g., flagging, categorizing, or filtering user submissions (Kumar et al. 2024). In academia, researchers may use LLMs to improve their manuscripts, while conferences and journals are beginning to experiment with LLM-assisted peer review (Thakkar et al. 2025). Similar patterns are emerging in education and customer service, where LLMs support both communication and assessment tasks (Forbes 2024). In these domains and beyond, LLMs increasingly both generate and evaluate the same content, giving rise to a new class of AI–AI interactions with significant implications for human decision-making and business operations (Laurito et al. 2025).
@@
Recent research in computer science has identified a behavioral tendency in LLMs known as self-preference, i.e., the inclination of a model to favor content it generated itself over that written by humans or produced by alternative models (Panickssery et al. 2024). While this phenomenon has been documented in benchmark evaluations (Zheng et al. 2023, Bai et al. 2023), its implications for real-world decision-making remain largely unexplored. As LLMs are increasingly deployed in high-stakes settings where they may evaluate content also generated by LLMs, self-preference introduces a novel form of bias. Unlike traditional biases rooted in demographic disparities (Sheng et al. 2019), this bias emerges endogenously from AI-AI interactions, in which the model’s own evaluative behavior systematically favors outputs aligned with its generative patterns.
@@
Self-preferencing bias is inherently interactional: it arises when LLMs are asked to judge content that may share stylistic or linguistic patterns with their own generative outputs. As such, it poses a new challenge for AI fairness, one that is not addressed by existing safeguards focused on demographic disparities. If left unchecked, self-preference could subtly distort evaluative processes across hiring, education, publishing, and more—privileging those who employ the same AI system used for evaluation (i.e., the “right” tool from the model’s perspective) while disadvantaging those who use different tools or none at all. Addressing this issue will require expanding current fairness frameworks to account for LLMs’ dual roles as both decision aids and evaluators in an increasingly AI-mediated business environment.
@@
Among the many domains where self-preference bias may arise, algorithmic hiring is particularly consequential. Employers are now routinely adopting LLMs to streamline resume screening and candidate ranking, often as part of automated workflows that support human decision-making (Gan et al. 2024, Kim et al. 2024, ResumeBuilder 2024, Sarumathi et al. 2025). Unlike traditional keyword-matching systems, LLMs can evaluate resumes in a more holistic manner—synthesizing content, inferring intent, and making contextual judgments beyond simple heuristics (Pritchett 2025). While this shift promises greater efficiency and scalability, it may also magnify the risk of bias if an LLM systematically favors resumes that reflect its own generative style. In such cases, evaluations may hinge less on the substantive quality of a candidate’s credentials and more by superficial stylistic alignment with the evaluator LLM, conferring unwarranted advantages on applicants who use the same model to compose their materials. In effect, this bias rewards access to specific generative technologies and penalizes those without it, even when applicants are otherwise equally qualified.
@@
In this paper, we provide the first empirical evidence that self-preference bias can distort candidate evaluations in algorithmic hiring. Specifically, we examine whether LLMs, when deployed as evaluators, systematically favor resumes they generated themselves over otherwise equivalent resumes written by humans or produced by alternative LLM models. To test this, we conduct a large-scale resume correspondence experiment using a real-world dataset of $2,245$ human-written resumes, sourced from a professional resume-building platform prior to the widespread adoption of generative AI. For each resume, we generate multiple counterfactual versions using a range of state-of-the-art LLMs, including GPT-4o, GPT-4o-mini, GPT-4-turbo, LLaMA 3.3-70B, Mistral-7B, Qwen 2.5-72B, and Deepseek-V3.111These models represent the state of the art among widely deployed LLMs at the time of our study. Although newer and more capable models will continue to emerge, the key pattern we examine—LLMs being used both to generate content and to evaluate that content—already appears in many application domains and is expected to remain common. Our focus is therefore on a structural feature of AI-mediated decision making rather than on the behavior of any specific model version. Having content quality controlled, we assess whether these LLMs exhibit systematic bias in favor of their own outputs when acting as evaluators.
@@
We distinguish between two forms of self-preference bias: LLM-vs-Human, where a model prefers its own generated content over a human-written equivalent; and LLM-vs-LLM, where a model favors its own output over content produced by an alternative (different) LLM. We find strong and consistent evidence of LLM-vs-Human self-preferencing across most models. Larger models, such as GPT-4o, GPT-4-turbo, DeepSeek-V3, Qwen-2.5-72B, and LLaMA 3.3-70B, exhibit particularly strong bias, exceeding $65\%$ even after controlling for content quality and reaching over $80\%$ for GPT-4o. By contrast, LLM-vs-LLM self-preferencing is more heterogeneous. DeepSeek-V3 shows the strongest bias in this setting, preferring its own outputs by $69\%$ against LLaMA 3.3-70B and $28\%$ against GPT-4o. GPT-4o and LLaMA 3.3-70B, in comparison, do not display consistent preferences when evaluating content generated by other models.
@@
To mitigate AI self-preference, we propose two simple yet effective strategies that directly target the underlying mechanism of self-recognition—a model’s ability to implicitly identify content it generated. The first strategy uses system prompting to explicitly instruct models to ignore the origin of resumes and focus only on substantive content. The second strategy employs a majority voting ensemble, combining the evaluator model with smaller models that exhibit weaker self-recognition, thereby diluting the bias of any single LLM. Across all tested LLMs, these interventions reduce LLM-versus-human self-preference by $17\%$ to $63\%$ in relative terms. These results demonstrate that while self-preference bias is widespread and consequential, it is not immutable: straightforward design interventions can substantially improve fairness in LLM-based hiring operations.
Together, these findings advance our understanding of how generative AI systems shape operational decision processes in organizations. We document and quantify self-preference bias in the context of resume screening by developing a measurement framework grounded in established fairness metrics. In addition, we show that self-preference bias can be substantially reduced through targeted interventions informed by LLM’s self-recognition behavior. In doing so, the study introduces a novel and practical perspective on AI fairness for business operations—one that moves beyond concerns about demographic disparities to address interactional biases that arise when AI systems evaluate content they themselves could have produced.
@@ 2 Literature Review
Our research contributes to multiple strands of literature: (1) screening and hiring as operational processes, (2) generative AI in labor market and workplace operations, (3) algorithmic fairness, bias, and responsible operations, and (4) LLMs as judges in decision making.
2.1 Screening and Hiring as Operational Processes
Operations Management (OM) research has long examined hiring and screening as core operational processes through which organizations allocate scarce evaluation capacity and form productive matches under uncertainty (Freeman 1983, Jovanovic 1979, Arlotto et al. 2014, Yoo et al. 2016, Koren 2026). This work positions workforce selection at the interface between operations and human resource management (Boudreau et al. 2003), conceptualizing recruitment and screening not as purely administrative functions but as operational processes through which firms transform applicant pools into productive labor inputs that support day-to-day operations.
A central implication of this lens is that hiring pipelines are typically multi-stage and capacity constrained. Upstream screening (e.g., resume review) funnels a large applicant pool into a limited set of downstream evaluations (e.g., interviews and assessments), where time, managerial attention, and specialist resources are scarce. As in other service and information-processing systems, errors at upstream triage have asymmetric and persistent consequences: false negatives irreversibly eliminate qualified candidates, while false positives consume scarce evaluation capacity and degrade downstream match quality. Recent work in platform and labor market operations further emphasizes how design choices in these screening stages shape congestion, equilibrium outcomes, and match efficiency. For example, Farajollahzadeh et al. (2025) show that interview-stage diversity interventions such as the Rooney Rule can backfire in capacity-constrained settings with noisy pre-interview signals, while Baek et al. (2025) demonstrate that strategic interview decisions in multistage, multi-firm hiring processes can substantially mitigate market inefficiencies and increase the number of hires overall.
Our paper contributes to this stream of literature by identifying a new and operationally consequential distortion that arises at the earliest stage of the hiring funnel when screening is delegated to AI evaluators. Existing OM models implicitly assume that evaluators process signals in a manner that is neutral to how those signals are generated. We relax this assumption and offer empirical evidence by studying an environment in which applicants increasingly use LLMs to generate screening inputs (e.g., resumes), while firms deploy similar models to evaluate those same inputs. We show that AI evaluators can exhibit systematic self-preference, favoring content generated by the same model, even when underlying candidate information is held constant. This interaction-driven bias is distinct from standard screening noise: it is endogenous to pipeline design choices (which tools are used upstream and downstream) and can create path-dependent lock-in of “dominant” resume styles, with implications for both operational performance (misranking and misallocation) and responsible hiring system design.
2.2 Generative AI in Labor Market and Workplace Operations
A growing literature examines the role of AI in labor market and workplace operations. Research spanning operations management, economics, and information systems documents how AI tools are increasingly deployed in both production and evaluation tasks, with important implications for worker performance, learning, and market outcomes. For example, Brynjolfsson et al. (2025) provide field evidence that a generative AI assistant improves productivity in customer support, with heterogeneous effects depending on workers’ experience and baseline skill. In labor-market contexts, a complementary stream studies how AI-assisted writing affects the formation of applicant signals and subsequent hiring outcomes. Wiles et al. (2025) show that algorithmic resume writing assistance improves writing quality and increases hiring, suggesting that generative AI can meaningfully alter the informational content and effectiveness of applicant signals in matching markets. At the same time, the introduction of algorithms into evaluation processes raises well-documented concerns. Prior work shows that human decision makers may discount or resist algorithmic recommendations (Dietvorst et al. 2015), and that algorithmic hiring systems can encode or amplify existing biases when trained on historical data or used with discretion (Hoffman et al. 2018, Cowgill 2018).
Our paper contributes to this emerging body of work by identifying a distinct and previously unexamined mechanism that arises when generative AI is used on both sides of the hiring market. We provide controlled, large-scale evidence that such coupling can create systematic advantages for certain applicants even when quality is held constant. Specifically, we document and quantify AI self-preferencing in resume screening. We then connect this bias to labor-market consequences by simulating capacity-constrained shortlisting pipelines across occupations, showing how evaluator self-preference affects the allocation of interview opportunities. Finally, we demonstrate that simple interventions (e.g., system prompting and majority-vote ensemble) can substantially mitigate this distortion, offering practical guidance for firms adopting generative AI in hiring workflows.
2.3 Algorithmic Fairness, Bias, and Responsible Operations
Our work also contributes to the growing literature on responsible practices in operations management, which emphasizes that operational decisions should be evaluated not only by efficiency but also by their broader implications for fairness and stakeholder welfare. Recent research highlights that dimensions such as fairness, equity, and employee well-being are central—rather than peripheral—to the design of effective and sustainable operational systems (Lee and Tang 2018, Netessine 2022, Kesavan et al. 2022, Wang et al. 2023). This body of work underscores a core insight: responsible operations require careful attention to how system design choices shape downstream human outcomes, often in context-specific ways.
@@ 2.3 Algorithmic Fairness, Bias, and Responsible Operations
Hiring and screening provide a particularly salient application of these ideas. Discrimination in labor markets has long been a central concern for both policymakers and economists, with extensive research documenting that certain race, gender, and age subpopulations face unfair treatment from recruiters in the hiring process (Bertrand and Mullainathan 2004, Kline et al. 2022, Neumark 2018). Concerningly, studies such as Datta et al. (2015) and Lambrecht and Tucker (2019) have further shown that, beyond human bias, algorithms deployed in hiring systems can also encode discriminatory behavior that disadvantages particular demographic groups. In recent years, the rise of LLMs with their increasing adoption in hiring systems has prompted researchers to empirically evaluate whether emerging AI-driven hiring systems perpetuate similar forms of algorithmic discrimination. For instance, Veldanda et al. (2023) replicated Bertrand and Mullainathan’s correspondence experiment to investigate AI-driven hiring bias across gender, race, maternity status, pregnancy status, and political affiliation. Additionally, An et al. (2024) explored whether LLMs exhibit race- and gender-based discrimination through an experiment, in which LLMs write an email to a named job candidate about a hiring decision. Likewise, Nghiem et al. (2024) analyzed the extent to which LLMs exhibit bias toward applicants based on their first names when making employment recommendations.
At the same time, a broader interdisciplinary literature documents additional risks associated with the use of LLMs in organizational decision making, including hallucinations, systematic errors, and the reinforcement of social disparities (Magesh et al. 2024, Huang et al. 2025). These concerns have motivated extensive research on the ethical, legal, and governance implications of algorithmic hiring systems (Raghavan et al. 2020, Li et al. 2021, Hunkenschroer and Kriebitz 2023). For instance, Raghavan et al. (2020) study how firms that develop algorithmic pre-employment assessments build, validate, and manage bias in practice, highlighting gaps between technical capabilities, organizational incentives, and regulatory oversight.
Despite this growing attention to algorithmic fairness and governance, existing work largely evaluates hiring algorithms in isolation and focuses on disparities across protected attributes. Unlike prior work that centers on protected characteristics, the self-preference bias we study arises from interactions among algorithmic components within the hiring pipeline itself. By uncovering this interaction-induced bias, our study highlights an emergent risk in algorithmic hiring systems that is not directly addressed by existing fairness audits or regulatory frameworks.
More broadly, our findings inform ongoing debates on AI governance and responsible operations by demonstrating how seemingly neutral design choices, such as deploying the same or similar models for content generation and evaluation, can systematically distort how scarce evaluation capacity is allocated across candidates. This perspective underscores the need for regulatory and organizational approaches that account not only for input data and protected attributes, but also for the interaction structure of AI-driven operational processes.
2.4 LLMs as Judges in Decision Making